# Forecasting Local Ionospheric Parameters Using Transformers

### v1

@@ 3.1 Transformer Component
Our transformer is based on the original encoder-decoder architecture introduced by [38] which uses multi-head self-attention networks followed by fully connected feed-forward layers. The attention mechanism allows the transformer to discover long-range temporal relationships very efficiently and to incorporate information from the covariate time series into its forecast of each quantile in Equation 10. This is a desirable property for ionospheric forecasting, since there may be unknown time lags between changes in global driver indices and their effects on the target parameters [18, 19, 20, 21, 22].
@@
After each attention layer, we apply a fully connected feed-forward network with a nonlinear activation. For our model, we use GELU activation functions [39]. Although the original transformer architecture used different hidden dimensions in the feed-forward layer, $d_{ff}$, than in the attention mechanisms, we found that our model performed well with $d_{ff}=d_{m}=128$. These feed-forward networks introduce additional nonlinearity to the transformer’s processing of the data through their activation functions,
@@
where $\sigma(\cdot)$ is the nonlinear activation. We also incorporate normalization layers [40] after each attention and feed-forward step to improve the training process.
The output from Equations 27 and 28 are now ready for processing in the decoder’s cross-attention. Here, a new set of queries, keys, and values are created. The queries use output from the decoder’s self-attention, and the keys and values use output the encoder’s self-attention,
@@
The decoder’s cross-attention layer performs the critical function of fusing the information from the encoder with that of the decoder. Larger transformer models may include several blocks of encoders and decoders, each of which computes a new set of attention heads and repeats the entire process outlined above. These models will have many more parameters to train and thus require very large amounts of data, however, they offer more flexibility by allowing each new set of attention heads in each layer to attend to different information. For our model, we found that a single block of attention heads in both our encoder and decoder performed well and kept the total number of trainable parameters much lower than other transformer implementations for time series forecasting [41, 25]. Figure 7 provides a visualization of the attention scores for each of the model’s layers for an encoder input length of $c=288$ and a decoder input length of $p=96$. This corresponds to a context length of 3 days and a prediction length of 1 day.
@@
A final note is that, in practice, transformers often utilize residual connections, as introduced by [42], which simply add the input and output of a given layer together before being passed into the following layer, e.g. Equation 27 would be implemented as
@@
This is now common practice in machine learning models that involve many layers of neural networks, as residual connections have been known to help alleviate vanishing gradients and improve overall stability during training [43]. However, because our model consists of a single block of attention layers, i.e. one pass through the encoder and one pass through the decoder, these skip connections are not required. Nevertheless, we include them here since future efforts in the space weather community may require more expressive models, and one way of achieving this is by including additional blocks.
@@ 3.2 Loss Function
This loss function penalizes overestimation and underestimation asymmetrically based on the level of $\tau$, and when $\tau=0.5$ it is equivalent to the mean absolute error, or L1 loss function. This loss has been used as a standard forecasting metric for weather models where observations can only be obtained once [44, 45, 46, 47]. Figure 8 plots this loss function for several quantile levels, which serves to illustrate the asymmetric penalty induced by quantile levels above or below $\tau=0.5$.
@@
Using our seven selected quantile levels, we may generate empirical estimates of the $90^{th}-$percent CI as well as the interquartile range (IQR), defined by the region between the $25^{th}$ and $75^{th}$ percentiles. Conceptually, our transformer network is adding a nonlinear corrective term to the linear predictor by learning to predict the median residual, while also quantifying uncertainty in the forecast in a manner that incorporates all available data from both past and future and without relying on any assumptions of the underlying distributions of the residuals.
In practice, we also dynamically weight each component of the loss function, as the loss can become dominated by the errors of the naïve linear predictor $\mathcal{L}_{\text{linear}}$. We balance the loss function using learnable weights from each model component, $w_{\text{linear}}$ and $w_{\text{residual}}$, so that
@@
These weights adjust the contribution of each loss component during training, and the exponentiation ensures that they are always positive and inversely related to the magnitude of the respective component [48]. We found that this weighting reduced the number of epochs needed to reach our early stopping criterion, which was set to halt training when the validation loss was not reduced by at least $1\text{e}^{-4}$ for more than 8 epochs. Additionally, we include dropout layers [49] in each of the transformer’s multi-head attention blocks to further control overfitting.
The linear error term only depends on the parameters of each linear model component and thus can be pre-trained using multiple least squares fits. However, this could introduce difficulties if the LIFT model is trained on larger datasets, as the least squares equations would require significant memory. We also found that including the linear components as additional linear layers and training them with Adam did not significantly increase training time.
@@ 3.3 Positional Encoding
The basic transformer architecture does not have any inherent mechanism to preserve the temporal or sequential structure of the data. Therefore, it is often necessary to provide a positional encoding feature. The now conventional positional encoding from [38] only gives relative position information for each data segment, so, we also equip our transformer with a time vector. The time vector provides the model with absolute position information, which we split into quadrature components,
[…]
where $\text{YOS}_{t}$, $\text{DOY}_{t}$, $\text{HOD}_{t}$, and $\text{MOH}_{t}$ are components of the time variable $t$ that represent the year-of-solar-cycle, day-of-year, hour-of-day, and minute-of-day respectively [50, 51].
@@ 3.4 Model covariates
Incorporating additional covariates into the transformer model is as easy as concatenating another time series to the matrix $\mathbf{X}$. Moreover, what sets the transformer apart from most other empirical and assimilative models is its natural ability to incorporate what may be considered future-known covariates through the decoder block (see Figure 6). Since IRI is a quick and efficient means of generating climatological predictions of the F2-layer parameters, it is an obvious choice for a decoder input. We use a convenient, pure Python version of IRI called PyIRI [52] to generate foF2, hmF2, and the thickness parameters $B_{bot}^{\text{F2}}$ and $B_{top}^{\text{F2}}$ for every contiguous data segment. Together, these IRI covariates allow the transformer to include climatology for the F2-region of the ionosphere that uses the latest parameterization of the observed F10.7cm solar flux. Finally, the solar zenith angle (SZA) is calculated for each geographic location and time [53]. Figure 9 illustrates these additional covariates for the Boulder, CO, USA ionosonde station.
Figure 9: Model covariates that form $\mathbf{Z}_{t-c:t+p}$, which are computable and provide the encoder and decoder networks with climatology (IRI) and solar position data (SZA) for a given sounding station.
@@ 4 Results
A challenge with any time series model is determining the appropriate length of data to provide as input. In classic Box-Jenkins methods, a simple partial autocorrelation can inform the maximum input sequence length, and while previous work [18] suggests that delayed responses of certain ionospheric parameters to solar activity can exceed 48 hours, giving a rough estimate for a desired time lag, our transformer is learning to embed all data in high-dimensional latent variables, thus potentially altering any correlations one may find through the raw data alone. To address this, we adopt a brute-force method and train multiple models with different sequence lengths for the encoder and decoder networks. This was only viable given the relatively reasonable training times of roughly 1-5 hours per model, which we carried out in parallel on a computer with 4 NVIDIA A100 GPUs. Figure 10 shows the result of this model search in which we find a clear relationship between longer encoder lengths and a reduction in the overall root mean squared error (RMSE).
@@
We are primarily interested in the model’s ability to produce a 24-hour forecast; Figure 11 shows the same RMSE values, restricted to a decoder length of 24 hours. In light of Figures 10 and 11, we chose an input sequence length of 3 days: beyond 3 days, we find that the total RMSE for a 24-hour forecast is only marginally improved with longer input sequences. This input length balances the accuracy of the forecast against the need for very long sequential segments from a given sounder, which makes the model more practical in an operational setting. In practice, of course, one may retrain this model to fit their specific needs for forecast horizon.
@@
The test stations, which were held out entirely during the training process, allow us to examine the model’s ability to generalize to new geographic locations. Randomly selected samples from the test set are provided in the following figures to illustrate the prediction of foF2, hmF2, and TEC in practice along with their respective $90\%$ CI and IQR. Figure 12 is from a mid-latitude sounder located at $(41.90\text{N},12.50\text{E})$, and additional forecasts for high-latitude $(64.66\text{N},212.93\text{E})$ and low-latitude $(-2.60\text{N},315.80\text{E})$ stations are given in Figures 13 and 14, respectively. These forecasts show, qualitatively, the model’s ability to produce uncertainty quantification that is dynamic and dependent on the given segment of data.
Figure 12: Model forecasts for foF2, hmF2, and TEC parameters from a mid-latitude test holdout station RO041 in Rome, Italy.
Figure 13: Model forecasts for foF2, hmF2, and TEC parameters from a high-latitude test station EI764 near Fairbanks, Alaska, USA.
Figure 14: Model forecasts for foF2, hmF2, and TEC parameters from a low-latitude test station SAA0K in Sao Luis, Brazil.
@@ 4.1 Forecast accuracy
We quantify the overall forecast accuracy of the combined linear model plus median predictions from the transformer using Figures 15, 16, and 17. These figures report the median absolute deviation (MAD) and mean absolute percent error (MAPE) in addition to RMSE. We achieve an overall RMSE of 0.617 MHz for foF2 across all test segments, including high- and low- solar activity and different latitudes. The hmF2 parameter, which is a challenging parameter to predict but has a high impact on many operational systems [5], has an average RMSE of 21.6 km, and the TEC parameter averages 2.426 TECU over the entire test period. Note that the IQRs for each are represented in the box-plots underneath each error distribution and show that the forecasts, in general, are capturing the point predictions well.
@@
Figure 15: Model forecast statistics for the foF2 parameter averaged across all test stations.
@@
Figure 16: Model forecast statistics for the hmF2 parameter averaged across all test stations.
@@
Figure 17: Model forecast statistics for the TEC parameter averaged across all test stations.
Comparisons with IRI are straightforward, since PyIRI has provided the peak height and frequency of the F2-layer as additional covariates to the LIFT encoder and decoder. Figure 18 shows the RMSE error distribution for foF2 and hmF2 using LIFT versus IRI. The LIFT model provides an appreciable increase in skill over IRI parameterized with the F10.7cm solar flux value from the last observation before the forecast period begins. Despite the best available drivers, IRI is still a climatological model, and without computationally expensive data assimilation schemes [54], the short-term forecast accuracy is difficult to improve.
@@
Figure 18: Model median predictions versus IRI parameterized with the F10.7 value from at the final context variable time step.
@@
The LIFT model provides uncertainty quantification with each forecast; we evaluate how well calibrated the predicted quantiles are through Figure 19, in which each distribution represents the percentage of points that fall within each confidence region. The statistics for these confidence regions are determined over the full 24-hour duration of each forecast individually,
[…]
IQR $\text{IQR Calibration}(\hat{\mathbf{x}}_{t+1:t+p;j})$ $=100*\frac{1}{p}\sum_{i=1}^{p}\mathbbm{1}(\hat{x}^{0.25}_{t+i;j}<x_{t+i;j}<\hat{x}^{0.75}_{t+i;j}),$
where $\mathbbm{1}(condition)$ is the indicator function which equals 1 when condition is satisfied and 00 otherwise. So, for each 24-hour forecast from LIFT, we obtain a single percentage for the number of points correctly within each CI range. Perfectly calibrated quantiles over the entire test dataset would result in the distributions in Figure 19 being delta functions at 90 and 50, for the 90% CI and IQR, respectively. Here we see that both the 90% CI and IQR are conservative, often estimating uncertainty bounds that are slightly too wide. Future iterations of this approach may explore additional means of calibrating the uncertainty bounds, such as incorporating conformalized quantiles [55], however, our results for the 90% CI for TEC are comparable to those of [56] who explored the use of super ensembles, quantile gradient boosting, and Bayesian neural network methods for uncertainty quantification.
@@
Figure 20: Boxplots of the daily percentile values for (top) foF2 (MHz), (middle) hmF2 (km), and (bottom) total electron content (TEC, in TECU). For each day in the test dataset, the specified percentile (e.g., 10%, 20%, …, 90%) is computed from the observed data (black), the forecast model predictions (colored boxes, as labeled), and the IRI climatological predictions (colored boxes, as labeled). Each box therefore represents the distribution across all days of that daily p-th percentile; the box edges show the interquartile range, the horizontal line (or triangle) indicates the median (or mean), and whiskers extend to capture the spread of values. Comparison of the model and IRI distributions against the observed distribution provides an overall indication of forecast bias and variability at different percentile levels.
We may also examine our overall model bias as compared to the IRI climatological predictions. Figure 20 demonstrates an improvement in the total model bias using LIFT over IRI. Each boxplot corresponds to the distribution across all 24‐hour forecasted periods in the test dataset of the p-th percentile for observed data, the forecast model, and the climatological model (IRI). That is, for each day we compute, say, the $10^{th}$ percentile of the observed data, and then gather those $10^{th}$‐percentile values across all days into one boxplot—repeating the same procedure for the model and for IRI at each percentile. From the figure, we show how the median, interquartile range, and whiskers for that daily percentile varies over the entire test dataset. While this demonstrates a generally well-calibrated model over the entirety of each 24-hour forecast, calibration of the model’s uncertainty as a function of time-of-day will be a focus for future efforts.

### v2

@@ 3.1 Transformer Component
Our transformer is based on the original encoder-decoder architecture introduced by (?, ?) which uses multi-head self-attention networks followed by fully connected feed-forward layers. The attention mechanism allows the transformer to discover long-range temporal relationships very efficiently and to incorporate information from the covariate time series into its forecast of each quantile in Equation 10. This is a desirable property for ionospheric forecasting, since there may be unknown time lags between changes in global driver indices and their effects on the target parameters (?, ?, ?, ?, ?, ?).
@@
After each attention layer, we apply a fully connected feed-forward network with a nonlinear activation. For our model, we use GELU activation functions (?, ?). This choice of activation function is simply because GELU provides modest training stability and convergence in other transformer applications; however, a ReLU function would likely work as well here. Although the original transformer architecture used different hidden dimensions in the feed-forward layer, $d_{ff}$, than in the attention mechanisms, we found that our model performed well with $d_{ff}=d_{m}=128$. These feed-forward networks introduce additional nonlinearity to the transformer’s processing of the data through their activation functions,
@@
where $\sigma(\cdot)$ is the nonlinear activation. We also incorporate normalization layers (?, ?) after each attention and feed-forward step to improve the training process.
The output from Equations 27 and 28 are now ready for processing in the decoder’s cross-attention. Here, a new set of queries, keys, and values are created. The queries use output from the decoder’s self-attention, and the keys and values use output from the encoder’s self-attention,
@@
The decoder’s cross-attention layer performs the critical function of fusing the information from the encoder with that of the decoder. Larger transformer models may include several blocks of encoders and decoders, each of which computes a new set of attention heads and repeats the entire process outlined above. These models will have many more parameters to train and thus require very large amounts of data; however, they offer more flexibility by allowing each new set of attention heads in each layer to attend to different information. For our model, we found that a single block of attention heads in both our encoder and decoder performed well and kept the total number of trainable parameters much lower than other transformer implementations for time series forecasting (?, ?, ?). Figure 8 provides a visualization of the attention scores for each of the model’s layers for an encoder input length of $c=288$ and a decoder input length of $p=96$. This corresponds to a context length of 3 days and a prediction length of 1 day.
@@
A final note is that, in practice, transformers often utilize residual connections, as introduced by (?, ?), which simply add the input and output of a given layer together before being passed into the following layer, e.g. Equation 27 would be implemented as
@@
This is now common practice in machine learning models that involve many layers of neural networks, as residual connections have been known to help alleviate vanishing gradients and improve overall stability during training (?, ?). However, because our model consists of a single block of attention layers, i.e. one pass through the encoder and one pass through the decoder, these skip connections are not required. Nevertheless, we include them here since future efforts in the space weather community may require more expressive models, and one way of achieving this is by including additional blocks. However, additional attention blocks will greatly increase the parameter count of the model.
@@ 3.2 Loss Function
This loss function penalizes overestimation and underestimation asymmetrically based on the level of $\tau$, and when $\tau=0.5$ it is equivalent to the mean absolute error, or L1 loss function. This loss has been used as a standard forecasting metric for weather models where observations can only be obtained once (?, ?, ?, ?, ?). Figure 9 plots this loss function for several quantile levels, which serves to illustrate the asymmetric penalty induced by quantile levels above or below $\tau=0.5$.
@@
Using our seven selected quantile levels, we may generate empirical estimates of the $90^{th}$-percent CI as well as the interquartile range (IQR), defined by the region between the $25^{th}$ and $75^{th}$ percentiles. We explored the use of different quantiles, i.e. $\tau={0.01,0.02,\dots,0.99}$, but found that specifying more quantiles did not increase the calibration significantly. The choice here was based only on the desired uncertainty quantification regions. Additionally, using fewer quantiles slightly reduces the number of model parameters and thus increases training speed. This is due to the smaller output dimension of the final feedforward layers in the decoder, Equations 34, 35, and 36.
Conceptually, our transformer network is adding a nonlinear corrective term to the linear predictor by learning to predict the median residual, while also quantifying uncertainty in the forecast in a manner that incorporates all available data from both past and future. It does this without relying on any assumptions of the underlying distributions of the residuals. In practice, we also dynamically weight each component of the loss function, as the loss can become dominated by the errors of the naïve linear predictor $\mathcal{L}_{\text{linear}}$. We balance the loss function using learnable weights from each model component, $w_{\text{linear}}$ and $w_{\text{residual}}$, so that
@@
These weights adjust the contribution of each loss component during training, and the exponentiation ensures that they are always positive and inversely related to the magnitude of the respective component (?, ?). We found that this weighting reduced the number of epochs needed to reach our early stopping criterion, which was set to halt training when the validation loss was not reduced by at least $1\text{e}^{-4}$ for more than 8 epochs. Additionally, we include dropout layers (?, ?) in each of the transformer’s multi-head attention blocks to further control overfitting.
The linear error term only depends on the parameters of each linear model component and thus can be pre-trained using multiple least squares fits. However, this could introduce difficulties if the LIFT model is trained on larger datasets, as the least squares equations would require significant memory. We also found that including the linear components as additional linear layers and training them with Adam did not significantly increase training time. To train the linear and residual components more effectively, we detach the gradients for the linear loss term from the residual loss term. This stops the quantile loss from updating the linear head. It makes the residual used by the quantile term a constant with respect to the linear head, so the linear head is trained only by MSE and the quantile head is trained to model the distribution of those residuals
@@ 3.3 Positional Encoding
The basic transformer architecture does not have any inherent mechanism to preserve the temporal or sequential structure of the data. Therefore, it is often necessary to provide a positional encoding feature. The now conventional positional encoding from (?, ?) only gives relative position information for each data segment, so, we also equip our transformer with a time vector. The time vector provides the model with absolute position information, which we split into quadrature components,
$\mathbf{t}(t)$ $=\begin{bmatrix}\cos\left(\frac{2\pi\text{YOS}_{t}}{11\times 365.25}\right)\\ \cos\left(\frac{2\pi\text{DOY}_{t}}{365.25}\right)\\ \cos\left(\frac{2\pi\text{HOD}_{t}}{24}\right)\\ \cos\left(\frac{2\pi\text{MOH}_{t}}{60}\right)\\ \sin\left(\frac{2\pi\text{YOS}_{t}}{11\times 365.25}\right)\\ \sin\left(\frac{2\pi\text{DOY}_{t}}{365.25}\right)\\ \sin\left(\frac{2\pi\text{HOD}_{t}}{24}\right)\\ \sin\left(\frac{2\pi\text{MOH}_{t}}{60}\right)\end{bmatrix}^{T},$ (42)
where $\text{YOS}_{t}$, $\text{DOY}_{t}$, $\text{HOD}_{t}$, and $\text{MOH}_{t}$ are components of the time variable $t$ that represent the year-of-solar-cycle, day-of-year, hour-of-day, and minute-of-hour respectively (?, ?, ?).
@@ 3.4 Model covariates
Incorporating additional covariates into the transformer model is as easy as concatenating another time series to the matrix $\mathbf{X}$. Moreover, what sets the transformer apart from most other empirical and assimilative models is its natural ability to incorporate what may be considered future-known covariates through the decoder block (see Figure 7). Since PyIRI is a quick and efficient means of generating climatological predictions of the F2-layer parameters, it is an obvious choice for a decoder input. We use PyIRI to generate foF2, hmF2, and the thickness parameters $B_{bot}^{\text{F2}}$ and $B_{top}^{\text{F2}}$ for every contiguous data segment. Together, these PyIRI covariates allow the transformer to include climatology for the F2-region of the ionosphere that uses the latest parameterization of the observed F10.7 cm solar flux. Finally, the solar zenith angle (SZA) is calculated for each geographic location and time (?, ?). Figure 10 illustrates these additional covariates for the Boulder, CO, USA ionosonde station.
Figure 10: Model covariates that form $\mathbf{Z}_{t-c:t+p}$, which are computable and provide the encoder and decoder networks with climatology (PyIRI) and solar position data (SZA) for a given sounding station.
@@ 4 Results
A challenge with any time series model is determining the appropriate length of data to provide as input. In classic Box-Jenkins methods, a simple partial autocorrelation can inform the maximum input sequence length, and while previous work (?, ?) suggests that delayed responses of certain ionospheric parameters to solar activity can exceed 48 hours, giving a rough estimate for a desired time lag, our transformer is learning to embed all data in high-dimensional latent variables, thus potentially altering any correlations one may find through the raw data alone. To address this, we adopt a brute-force method and train multiple models with different sequence lengths for the encoder and decoder networks. This was only viable given the relatively reasonable training times of roughly 1-5 hours per model, which we carried out in parallel on a computer with 4 NVIDIA A100 GPUs. Figure 11 shows the result of this model search in which we find a clear relationship between longer encoder lengths and a reduction in the overall root mean squared error (RMSE).
@@
We are primarily interested in the model’s ability to produce a 24-hour forecast; Figure 12 shows the same RMSE values, restricted to a decoder length of 24 hours. In light of Figures 11 and 12, we chose an input sequence length of 3 days: beyond 3 days, we find that the total RMSE for a 24-hour forecast is only marginally improved with longer input sequences. This input length balances the accuracy of the forecast against the need for very long sequential segments from a given sounder, which makes the model more practical in an operational setting. In practice, of course, one may retrain this model to fit their specific needs for forecast horizon.
@@
The test stations, which were held out entirely during the training process, allow us to examine the model’s ability to generalize to new geographic locations. Randomly selected samples from the test set are provided in the following figures to illustrate the prediction of foF2, hmF2, and TEC in practice along with their respective $90\%$ CI and IQR. Figure 13 is from a mid-latitude sounder located at $(41.90\text{N},12.50\text{E})$, and additional forecasts for high-latitude $(64.66\text{N},212.93\text{E})$ and low-latitude $(-2.60\text{N},315.80\text{E})$ stations are given in Figures 14 and 15, respectively. These forecasts show, qualitatively, the model’s ability to produce uncertainty quantification that is dynamic and dependent on the given segment of data.
Figure 13: Example 24-hour LIFT forecast for a mid-latitude test station RO041 in Rome, Italy (geographic: $41.90^{\circ}$N, $12.50^{\circ}$E). The left portion of each panel shows the 72-hour input context (past observations), and the right portion shows the 24-hour forecast horizon. The solid red line indicates the LIFT median prediction, shaded regions represent the 90% confidence interval (blue) and interquartile range (green), and the yellow line shows the PyIRI climatological prediction.
Figure 14: Example 24-hour LIFT forecast for a high-latitude test station EI764 near Fairbanks, Alaska, USA (geographic: $64.66^{\circ}$N, $212.93^{\circ}$E). Format follows Figure 13.
Figure 15: Example 24-hour LIFT forecast for a low-latitude test station SAA0K in Sao Luis, Brazil (geographic: $2.60^{\circ}$S, $315.80^{\circ}$E). Format follows Figure 13.
@@ 4.1 Forecast accuracy
We quantify the overall forecast accuracy of the combined linear model plus median predictions from the transformer using Figures 16, 17, and 18. These figures report the median absolute deviation (MAD) and mean absolute percent error (MAPE) in addition to RMSE. We achieve an overall RMSE of 0.600 MHz for foF2 across all test segments, including high- and low- solar activity and different latitudes. The hmF2 parameter, which is a challenging parameter to predict but has a high impact on many operational systems (?, ?), has an average RMSE of 21.810 km, and the TEC parameter averages 2.385 TECU over the entire test period. Note that the IQRs for each are represented in the box-plots underneath each error distribution and show that the forecasts, in general, are capturing the point predictions well. These results are consistent with the performance reported in recent machine learning studies: 24-hour TEC forecasts typically achieve RMSE in the range of 2–3 TECU (?, ?, ?), and Informer-based foF2 models report RMSE values of 0.5–0.8 MHz depending on forecast horizon and station characteristics (?, ?, ?). Direct numerical comparison across studies is complicated by differences in test station selection, temporal coverage, and evaluation protocols; nevertheless, LIFT’s performance falls within the range of current state-of-the-art methods while additionally providing joint multi-parameter forecasts with calibrated uncertainty quantification.
@@
Figure 16: Model forecast RMSE for the foF2 parameter averaged across all forecasted time periods.
@@
Figure 17: Model forecast RMSE for the hmF2 parameter averaged across all forecasted time periods.
@@
Figure 18: Model forecast RMSE for the TEC parameter averaged across all forecasted time periods.
We compare LIFT against PyIRI to quantify the skill improvement achievable by incorporating recent ionospheric observations over climatology alone. Comparisons against IRI-family models are standard practice in the ionospheric forecasting literature, including recent machine learning studies (?, ?, ?, ?), and PyIRI provides a reproducible, well-understood baseline. Direct comparisons with other machine learning models are difficult to conduct in a controlled manner, as most existing transformer-based ionospheric models focus on global TEC maps rather than local ionosonde-derived parameters, train on substantially shorter data records, and do not jointly forecast foF2, hmF2, and TEC with uncertainty quantification (?, ?). Comparisons with PyIRI are also straightforward to implement, since it provides the peak height and frequency of the F2-layer as additional covariates to the LIFT encoder and decoder. Figure 19 shows the RMSE error distribution for foF2, hmF2, and TEC using LIFT versus PyIRI. The LIFT model provides an appreciable increase in skill over PyIRI parameterized with the F10.7cm solar flux value from the last observation before the forecast period begins. We also perform a permutation test and generate bootstrap CIs on each set of distributions in Figure 19 to ensure we have statistical significance in the difference of means. P-values and confidence intervals can be see in Table 2, all of which show that LIFT’s RMSE is statistically significantly lower than PyIRI at an $\alpha=0.05$ level. Despite the best available drivers, PyIRI is still a climatological model, and without computationally expensive data assimilation schemes (?, ?), the short-term forecast accuracy is difficult to improve. We also include Table 3, which provides a summary of the RMSE performance for the LIFT model as a comparison to a linear-only version of LIFT, a transformer-only version, and the PyIRI forecasts. The illustrates how the transformer and linear baseline complement each other and obtain lower RMSE.
@@
Figure 19: Model median predictions versus PyIRI parameterized with the F10.7 value at the final context variable time step.
Statistical Significance (RMSE mean distributions)
Forecast Target Perm. P-value Bootstrap 95% CI
foF2 (MHz) p=0.0004 [-0.745, -0.737]
hmF2 (km) p=0.0004 [-11.543, -11.320]
TEC (TECU) p=0.0004 [-3.532, -3.485]
Table 2: Statistical significance for the RMSE predictions for all test data. Permutation test P-values and bootstrap confidence intervals are reported for each of the target forecast variables. Null hypothesis is that the means of the RMSE distributions for LIFT and PyIRI are the same.
RMSE
Forecast Target LIFT Transformer only Linear only PyIRI
foF2 (MHz) 0.600 0.898 0.637 1.341
hmF2 (km) 21.810 32.111 22.487 33.241
TEC (TECU) 2.385 2.808 2.482 5.894
Table 3: Summary statistics for the RMSE for all test data. The RMSE is averaged over all points in the test dataset. Comparisons of LIFT to the transformer only and linear only components along with PyIRI. Best performance in RMSE for each target variable is indicated in bold.
@@
The LIFT model provides uncertainty quantification with each forecast. We evaluate how well calibrated the predicted quantiles are through Figure 20, in which each distribution represents the percentage of points that fall within each confidence region. The statistics for these confidence regions are determined over the full 24-hour duration of each forecast individually,
$\text{90\% CI Calibration}(\hat{\mathbf{x}}_{t+1:t+p;j})$ $=100*\frac{1}{p}\sum_{i=1}^{p}\mathbbm{1}(\hat{x}^{0.05}_{t+i;j}<x_{t+i;j}<\hat{x}^{0.95}_{t+i;j})$
$\text{IQR Calibration}(\hat{\mathbf{x}}_{t+1:t+p;j})$ $=100*\frac{1}{p}\sum_{i=1}^{p}\mathbbm{1}(\hat{x}^{0.25}_{t+i;j}<x_{t+i;j}<\hat{x}^{0.75}_{t+i;j}),$
where $\mathbbm{1}(condition)$ is the indicator function which equals 1 when condition is satisfied and 0 otherwise. So, for each 24-hour forecast from LIFT, we obtain a single percentage for the number of points correctly within each CI range. Perfectly calibrated quantiles over the entire test dataset would result in the distributions in Figure 20 being delta functions at 90 and 50, for the 90% CI and IQR, respectively. Here we see that both the 90% CI and IQR are quite well calibrated in terms of their overall mean performance; however, the spread of each distribution implies that uncertainty quantification of LIFT is often too large or too small for many test segments. Table 4 provides a summary of the LIFT model performance at different geomagnetic latitude bands. From this table, we see that the LIFT model maintains reasonably well calibrated uncertainty bands even at the low and high latitudes. We include full histograms of the LIFT quantile calibration stratified by geomagnetic latitude in the Supporting Information. Note, that there are far fewer segments of data to test on at the high latitudes due to the relatively sparse number of GIRO station locations in these regions. This is a limitation of the data available.
Future iterations of this approach should explore means of better calibrating the uncertainty bounds. Some recent work in this area includes incorporating conformalized quantiles (?, ?), quantile gradient boosting, and Bayesian neural network. While further improvements to the uncertainty calibration are left to future work, our results demonstrate a significant advantage over point forecasts by using a simple linear model for the point predictions and a quantile loss function for the residuals.
@@
Num. Test
Segments
RMSE
Percent Within
90% CI
Percent Within
IQR
Full Test Set 76,894
foF2: 0.600
hmF2: 21.810
TEC: 2.385
foF2: 89.047
hmF2: 87.952
TEC: 88.050
foF2: 48.317
hmF2: 47.634
TEC: 47.116
Low Geomag Lat
+/-($0^{\circ}$, $20^{\circ}$)
24,152
foF2: 0.760
hmF2: 26.466
TEC: 3.491
foF2: 87.620
hmF2: 88.062
TEC: 87.115
foF2: 46.727
hmF2: 47.833
TEC: 46.386
Mid Geomag Lat
+/-($20^{\circ}$, $60^{\circ}$)
52,598
foF2: 0.528
hmF2: 19.687
TEC: 1.880
foF2: 89.691
hmF2: 87.893
TEC: 88.467
foF2: 49.033
hmF2: 47.537
TEC: 47.441
High Geomag Lat
+/-($60^{\circ}$, $90^{\circ}$)
144
foF2: 0.319
hmF2: 16.267
TEC: 1.260
foF2: 93.229
hmF2: 90.676
TEC: 92.535
foF2: 53.328
hmF2: 50.014
TEC: 50.666
Table 4: Summary of the RMSE and confidence interval (CI) calibration of the LIFT model stratified across low, medium, and high geomagnetic latitude bands. For the $90^{th}$ percentile and the IQR uncertainty bands, perfectly calibrated forecasts would result in 90 and 50 respectively. See Supporting Information for full histograms of the low, mid, and high latitude stratified data.
Figure 21: Boxplots of the daily percentile values for (top) foF2 (MHz), (middle) hmF2 (km), and (bottom) total electron content (TEC, in TECU). For each day in the test dataset, the specified percentile (e.g., 10%, 20%, …, 90%) is computed from the observed data (black), the forecast model predictions (colored boxes, as labeled), and the PyIRI climatological predictions (colored boxes, as labeled). Each box therefore represents the distribution across all days of that daily p-th percentile; the box edges show the interquartile range, the horizontal line (or triangle) indicates the median (or mean), and whiskers extend to capture the spread of values. Comparison of the model and PyIRI distributions against the observed distribution provides an overall indication of forecast bias and variability at different percentile levels.
We may also examine our overall model bias as compared to the PyIRI climatological predictions. Figure 21 demonstrates an improvement in the total model bias using LIFT over PyIRI. Each boxplot corresponds to the distribution across all 24‐hour forecasted periods in the test dataset of the p-th percentile for observed data, the forecast model, and the climatological model. That is, for each day we compute, say, the $10^{th}$ percentile of the observed data, and then gather those $10^{th}$‐percentile values across all days into one boxplot—repeating the same procedure for the model and for PyIRI at each percentile. From the figure, we show how the median, interquartile range, and whiskers for that daily percentile varies over the entire test dataset. While this demonstrates a generally well-calibrated model over the entirety of each 24-hour forecast, further calibration of the model’s uncertainty as a function of time-of-day will be a focus for future efforts.
The close visual agreement between model and observed boxplots in Figure 21, particularly for TEC, reflects successful spatial generalization rather than overfitting. These results are derived from the 76,894 test segments shown in Table 4, which represent stations completely excluded from training and validation. The distributional alignment demonstrates that LIFT reproduces the statistical properties of ionospheric variability at novel locations, while individual forecasts retain quantifiable error as shown in the RMSE distributions of Figures 1618. We note that TEC, being derived from the integrated electron density profile, exhibits a more statistically regular distribution with fewer extreme outliers compared to foF2 and hmF2, which may contribute to the particularly close distributional agreement observed for this parameter.